# Well quasi-order and atomicity for combinatorial structures under consecutive orders

### v3

@@ 1. Introduction
We will study two order-theoretic properties for these posets: well quasi-order, often shortened to wqo (which corresponds to the absence of infinite antichains), and atomicity (the property of being indecomposable as the union of two proper downward closed subsets). We present our findings within the framework of two decidability questions:
@@ 2. Preliminaries
  • On the other hand, the $\mathcal{W}_{A}$ is not wqo whenever $|A|>1$ – for example, if $A=\{a,b\}$ then $aa,aba,abba,abbba,\dots$ forms an infinite antichain.
@@
Atomic sets are also known as ideals, ages and directed sets (for instance, in [8]). The following proposition gives an equivalent condition to atomicicty, which we will generally use rather than the previous definition; a proof can be found in [8, Section 2.3.11].
@@
The avoidance set $\operatorname{Av}(ab,ba)$ of $\mathcal{W}_{\{a,b\}}$ consists of all of the words containing only one letter. It is not atomic because it does not satisfy the JEP – for instance, there is no word in $\operatorname{Av}(ab,ba)$ containing both $aa$ and $bb$ as subwords. Indeed, it can be seen that $\operatorname{Av}(ab,ba)=\operatorname{Av}(a)\cup\operatorname{Av}(b)$.
@@ 3. Definitions and results from graph theory
A path in a digraph is a sequence $v_{1}\rightarrow\dots\rightarrow v_{n}$ of vertices; the number of edges n1n-1 is the path’s length. A path is simple if all of its vertices are distinct and is a cycle if $v_{1}=v_{n}$ and it has length $\geq 1$. A cycle is a simple cycle if its only repeated vertex is the start/end vertex.
@@ 4. Factor Graphs
We will now associate the avoidance sets of our posets with certain digraphs, which will enable us to encode structures as paths in these digraphs. This will lead to an investigation of the relationship between paths in the digraphs and structures in our avoidance sets, utilising observations such as Propositions 3.12 and 3.13.
In what follows, we consider an avoidance set $C=\operatorname{Av}(B)$ and let $b$ be the maximum length of an element of $B$. We will let $C_{Y}=\{\sigma\in C:|\sigma|\in Y\}$ and shorten $C_{\{m\}}$ to $C_{m}$.
@@ 5. Valid structures
First consider graphs. Let $G,H$ be two graphs on $p,q$ points respectively, and suppose $G\hskip-0.85358pt\!\upharpoonright_{[p-x+1,\hskip 1.99168ptp]}\cong H\hskip-0.85358pt\!\upharpoonright_{[1,x]}$ for some $x\geq 1$. If we identify the last $x$ points of $G$ with the first $x$ points of $H$, we obtain another graph, so graphs are valid structures.
@@
For linear orders: since of course linear orders contain a linear order naturally in their signatures, we do not need to add an additional linear order. Hence the only linear orders are of the form $1\leq 2\leq\dots\leq n$. Let $l_{1}=1\leq 2\leq\dots\leq n$ and $l_{2}=1\leq 2\leq\dots\leq m$ be two such linear orders, with $n\leq m$. It is easy to see that $l_{1}$ and $l_{2}$ overlap on $n$ points, and by identifying these points we obtain another copy of $l_{2}$. Hence linear orders are valid.
@@
For transitivity, let $a,b,c\in P\cup Q$ such that $a\leq_{pq}b$ and $b\leq_{pq}c$. If $a,b,c\in P$ (or Q), then transitivity is inherited from P (or Q). So now consider the case where $a,b\in P\backslash Q$ and $c\in Q\backslash P$. Here, $a\leq_{p}b$ and there exists $d\in P\cap Q$ such that $b\leq_{p}d\leq_{q}c$. Therefore, $a\leq_{p}b\leq_{p}d\leq_{q}c$, meaning that $a\leq_{pq}c$. The case where $a\in P\backslash Q$ and $b,c\in Q\backslash P$ is proved analogously. Next, consider the case where $a,c\in P\backslash Q$ and $b\in Q\backslash P$. Since $a\leq_{pq}b$, there exists $d\in P\cap Q$ such that $a\leq_{p}d\leq_{q}b$. Similarly, there exists $e\in P\cap Q$ such that $b\leq_{q}e\leq_{p}c$. We observe that $d\leq_{q}b\leq_{q}e$, so $d\leq_{q}e$ by transitivity of $\leq_{q}$. Since P and Q are identical on the points of $P\cap Q$, it follows that $d\leq_{p}e$ as well. Therefore, $a\leq_{p}d\leq_{p}e\leq_{p}c$, and so $a\leq_{p}c$ by transitivity of P. It follows that $a\leq_{pq}c$. The final case is proven similarly.
@@
Consider an avoidance set $C=\operatorname{Av}(B)$ of $(X,\leq)$ and let $m\geq b$. Take a path $\pi$ in $\Gamma_{C}^{m}$. We will use induction on the length of $\pi$. Suppose the length of $\pi$ is 0, say $\pi=v_{0}$. Since $v_{0}$ is a vertex, it is an element of $C$ by definition, so $v_{0}\in\Sigma(\pi)$ and $|\Sigma(\pi)|\geq 1$. Now suppose the length of $\pi$ is 1, say $\pi=v_{0}\rightarrow v_{1}$. This means that $v_{0}\hskip-0.85358pt\!\upharpoonright_{[2,\hskip 1.99168ptb]}\cong v_{1}\hskip-0.85358pt\!\upharpoonright_{[1,\hskip 1.99168ptb-1]}$, and so by the definition of validity there exists a structure $v$ on $b+1$ points such that $v\hskip-0.85358pt\!\upharpoonright_{[1,\hskip 1.99168ptb]}\cong v_{0}$ and $v\hskip-0.85358pt\!\upharpoonright_{[2,\hskip 1.99168ptb+1]}\cong v_{1}$. This structure $v$ has associated path $\pi$, meaning that $|\Sigma(\pi)|\geq 1$. Note that $v\in C$ because it is of the right structure type and contains no forbidden substructures.
Now suppose all paths of length <k<k in $\Gamma_{C}^{m}$ have at least one associated structure. Let $\pi=v_{0}\rightarrow\dots\rightarrow v_{k}$ be a path of length $k$ in $\Gamma_{C}^{m}$. By assumption the paths $v_{0}\rightarrow\dots\rightarrow v_{k-1}$ and $v_{k-1}\rightarrow v_{k}$ each have at least one associated structure; call these $\rho$ and $\sigma$ respectively. Now $\rho$ and $\sigma$ overlap on the $m$ points of $v_{k}$, and so since the structures of this poset are valid there exists a structure $\theta$ which combines $\sigma$ and $\rho$. It is clear that $\Pi(\theta)=\pi$, so $\theta\in\operatorname{Av}(B)$ and $|\Sigma(\pi)|\geq 1$. Again note that $\theta\in C$ because it is of the right structure type and contains no forbidden substructures. ∎
@@
Consider the factor graph in Figure 3, which is the factor graph of the avoidance set $C$ of $\mathcal{F}$ which avoids all forests on three points except for those of the vertices. Within this factor graph, we let $A$ be the vertex on the left and $B$ the vertex on the right. The path $\pi=A\rightarrow B$ is associated with the graph shown in Figure 4, which is a cycle, so certainly not a forest, and therefore not in $C$. This means that $\Sigma(\pi)=\varnothing$ and so $\mathcal{F}$ is not valid by the converse of Lemma 5.4.
@@
Figure 4. The graph corresponding to the path $\pi$ in Example 5.5
@@
We note that structures which can be defined as downward closed subsets of valid structures are not necessarily valid themselves. An example of this is given by overlap free words. A word over an alphabet $A$ is overlap free if it does not contain any subword of the form xyxyxxyxyx for $x\in A,y\in A^{*}$. To see that overlap free words are not valid structures, consider the set of overlap free words over $A=\{a,b\}$. Take $u=aba$, $v=baba$; clearly $u,v\in A^{*}$. Moreover, $u\hskip-0.85358pt\!\upharpoonright_{[2,3]}=ba=v\hskip-0.85358pt\!\upharpoonright_{[1,2]}$. However, identifying these two letters yields the word ababaababa, which is not overlap free.
In this case we can use the methodology for words under the consecutive order to answer the wqo and atomicity problems for overlap free words under the consecutive order. Let $\mathcal{OA}_{A}$ be the poset of overlap free words over an alphabet A under the consecutive order. We consider an avoidance set $C=\operatorname{Av}(B)$ of $\mathcal{OA}_{A}$. Take $Y=\{xyxyx\hskip 5.69054pt|\hskip 5.69054ptx\in A,y\in A^{*}\}$. It can be seen that $C$ is equal to the avoidance set $\operatorname{Av}(B\cup Y)$ of $\mathcal{W}_{A}$. Hence wqo and atomicity for $C$ are governed by wqo and atomicity respectively for this avoidance set of $\mathcal{W}_{A}$.
@@ 6. Bountiful structures
Let $\sigma,\rho$ be valid structures of the same type of length $m$ such that $\sigma\hskip-0.85358pt\!\upharpoonright_{[2,m]}\cong\rho\hskip-0.85358pt\!\upharpoonright_{[1,m-1]}$. If for any such $\sigma$ and $\rho$ there are at least two structures $\theta_{1},\theta_{2}$ of the same type as $\sigma$ and $\rho$ of length $m+1$, such that $\theta_{i}\hskip-0.85358pt\!\upharpoonright_{[1,\hskip 1.99168ptm]}\cong\sigma$ and $\theta_{i}\hskip-0.85358pt\!\upharpoonright_{[2,\hskip 1.99168ptm+1]}\cong\rho$ for $i=1,2$, then we say this type of structure is bountiful.
@@
We begin by showing that $(1)$ holds. Given any path $\pi=\sigma\rightarrow\rho$ in $\Gamma^{m}_{C}$, $\sigma\hskip-0.85358pt\!\upharpoonright_{[2,\hskip 1.99168ptm]}\cong\rho\hskip-0.85358pt\!\upharpoonright_{[1,\hskip 1.99168ptm-1]}$ so there are two structures $\theta_{1},\theta_{2}\in C$ on $m+1$ points such that $\theta_{i}\hskip-0.85358pt\!\upharpoonright_{[1,\hskip 1.99168ptm]}\cong\sigma$ and $\theta_{i}\hskip-0.85358pt\!\upharpoonright_{[2,\hskip 1.99168ptm+1]}\cong\rho$, for $i=1,2$, by the assumption of bountifulness. Each of these structures is in $\Sigma(\pi)$, so $|\Sigma(\pi)|\geq 2$ as required.
Now we will show that $(1)$ implies $(2)$. Since all paths of length 1 are ambiguous, all paths containing a subpath of length 1 are ambiguous, so all paths of length $\geq 1$ are ambiguous.
@@
We will look at linear orders in a little more detail. As noted in the proof of Lemma 5.3, in our framework the only linear order on $b$ points is $1\leq 2\leq\dots\leq m$. Therefore, for linear orders, $m$-dimensional factor graphs consist of a single vertex of this form with a loop on it. The only structure associated with a path of length one in such a factor graph is $1\leq 2\leq\dots\leq m\leq m+1$, meaning that condition $(1)$ of Lemma 6.2 is broken. Hence linear orders are not bountiful.

### v4

@@ 1. Introduction
We will study two order-theoretic properties for avoidance sets of these posets: well quasi-order, often shortened to wqo (which corresponds to the absence of infinite antichains), and atomicity (the property of being indecomposable as the union of two proper downward closed subsets). We present our findings within the framework of two decidability questions:
@@ 2. Preliminaries
  • On the other hand, the poset $\mathcal{W}_{A}$ is not wqo whenever $|A|>1$ – for example, if $A=\{a,b\}$ then $aa,aba,abba,abbba,\dots$ forms an infinite antichain.
@@
Atomic sets are also known as ideals, ages and directed sets (for instance, in [9]). The following proposition gives an equivalent condition to atomicity, which we will generally use rather than the previous definition; a proof can be found in [9, Section 2.3.11].
@@
The avoidance set $\operatorname{Av}(ab,ba)$ of $\mathcal{W}_{\{a,b\}}$ consists of all of the words containing either only $a$’s or only $b$’s. It is not atomic because it does not satisfy the JEP – for instance, there is no word in $\operatorname{Av}(ab,ba)$ containing both $aa$ and $bb$ as subwords. Indeed, it can be seen that $\operatorname{Av}(ab,ba)=\operatorname{Av}(a)\cup\operatorname{Av}(b)$.
@@ 3. Definitions and results from graph theory
A path in a digraph is a sequence $v_{1}\rightarrow\dots\rightarrow v_{n}$ of vertices; the number of edges n1n-1 is the path’s length. A path is simple if all of its vertices are distinct and is a cycle if $v_{1}=v_{n}$ and it has length $\geq 1$. A loop is a cycle of length 1. A cycle is a simple cycle if its only repeated vertex is the start/end vertex.
@@ 4. Factor Graphs
We will now associate the finitely based avoidance sets of our posets with certain digraphs, which will enable us to encode structures as paths in these digraphs. This will lead to an investigation of the relationship between paths in the digraphs and structures in our avoidance sets, utilising observations such as Propositions 3.12 and 3.13.
Throughout this paper, we will take $C=\operatorname{Av}(B)$ to be a finitely based avoidance set and let $b$ be the maximum length of an element of $B$. We will let $C_{Y}=\{\sigma\in C:|\sigma|\in Y\}$ and shorten $C_{\{m\}}$ to $C_{m}$.
@@
We briefly note that, since factor graphs are digraphs, they can have loops.
@@ 5. Valid structures
We will always be working with posets whose underlying sets consist of structures of a single type. As such, if this type of structure is valid, we will abuse the terminology slightly by referring to these structures as valid structures.
@@
Indeed, we would like to thank one of the anonymous reviewers for highlighting the following close relationship between validity and the strong amalgamation property. Denote by $T$ a type of structure in which there is no interaction between the distinguished linear order and other relations in the signature, and denote by $T^{-}$ the type of structure obtained by removing the distinguished linear order from $T$. Any linear ordering of the underlying set of a structure of type $T^{-}$ produces a structure of type $T$. We will briefly show that $T$ is valid if and only if $T^{-}$ satisfies the strong amalgamation property.
First suppose that $T$ is a valid type of structure. Take $\sigma$ to embed in $\rho$ and $\tau$ as an amalgam of structures of type $T^{-}$. We can add linear orders to $\sigma,\rho$ and $\tau$ so that $\sigma$ embeds in the last $|\sigma|$ elements of $\rho$ and the first $|\sigma|$ elements of $\tau$. The resulting structures are of type $T$ and so $\rho$ and $\tau$ can be $|\sigma|$-combined by the assumption of validity. Removing the linear order from the resulting structure yields a structure of type $T^{-}$ which witnesses that $T^{-}$ satisfies the strong amalgamation property.
For the reverse direction, suppose that $T^{-}$ satisfies the strong amalgamation property and take structures $\sigma,\rho$ of type $T$ which $x$-overlap for some $x$. Take $\theta$ to be the restriction of $\sigma$ (and $\rho$) to the points in this overlap. Let $\sigma^{-},\rho^{-},\theta^{-}$ be the structures of type $T^{-}$ obtained by removing the distinguished linear orders from $\sigma,\rho$ and $\theta$. Note that $\theta^{-}$ embeds in $\sigma^{-}$ and $\rho^{-}$. Since $T^{-}$ satisfies the strong amalgamation property, there is a structure $\tau^{-}$ such that both $\sigma^{-}$ and $\rho^{-}$ embed in $\tau^{-}$, the images of the copies of $\theta^{-}$ in $\sigma^{-}$ and $\rho^{-}$ coincide, and the intersection of the copies of $\sigma^{-}$ and $\rho^{-}$ in $\tau^{-}$ is precisely this copy of $\theta^{-}$. We can add a linear order to $\tau^{-}$ in such a way as to produce a structure $\tau$ of type $T$ which $x$-combines $\sigma$ and $\rho$, yielding validity of $T$.
@@
First consider graphs. Let $G,H$ be two graphs on $p,q$ points respectively, and suppose $G\hskip-0.85358pt\!\upharpoonright_{[p-x+1,\hskip 1.99168ptp]}\cong H\hskip-0.85358pt\!\upharpoonright_{[1,x]}$ for some $x\geq 1$. Identifying the last $x$ points of $G$ with the first $x$ points of $H$ uniquely determines the relative values of vertices with respect to the underlying linear order that governs consecutivity. Doing this, we obtain another graph, so graphs are valid structures.
@@
For linear orders: since of course linear orders contain a linear order naturally in their signatures, we do not need to add an additional linear order. Hence the only linear orders are of the form $1\leq 2\leq\dots\leq n$. Let $l_{1}=1\leq 2\leq\dots\leq n$ and $l_{2}=1\leq 2\leq\dots\leq m$ be two such linear orders, with $n\leq m$. It can be seen that $l_{1}$ and $l_{2}$ overlap on $x$ points for any $x\leq n$, and that these are the only possible overlaps. In any of these cases, by identifying the overlapping points we obtain another linear order. Hence linear orders are valid.
@@
For transitivity, let $a,b,c\in P\cup Q$ such that $a\leq_{pq}b$ and $b\leq_{pq}c$. If $a,b,c\in P$ (or Q), then transitivity is inherited from P (or Q). So now consider the case where $a,b\in P\backslash Q$ and $c\in Q\backslash P$. Here, $a\leq_{p}b$ and there exists $d\in P\cap Q$ such that $b\leq_{p}d\leq_{q}c$. Therefore, $a\leq_{p}b\leq_{p}d\leq_{q}c$, meaning that $a\leq_{pq}c$. Next, consider the case where $a,c\in P\backslash Q$ and $b\in Q\backslash P$. Since $a\leq_{pq}b$, there exists $d\in P\cap Q$ such that $a\leq_{p}d\leq_{q}b$. Similarly, there exists $e\in P\cap Q$ such that $b\leq_{q}e\leq_{p}c$. We observe that $d\leq_{q}b\leq_{q}e$, so $d\leq_{q}e$ by transitivity of $\leq_{q}$. Since P and Q are identical on the points of $P\cap Q$, it follows that $d\leq_{p}e$ as well. Therefore, $a\leq_{p}d\leq_{p}e\leq_{p}c$, and so $a\leq_{p}c$ by transitivity of P. It follows that $a\leq_{pq}c$. The other cases are proven similarly.
@@
Consider an avoidance set $C=\operatorname{Av}(B)$ of $(X,\leq)$ and let $m\geq b$. Take a path $\pi$ in $\Gamma_{C}^{m}$. We will use induction on the length of $\pi$. Suppose the length of $\pi$ is 0, say $\pi=v_{0}$. Since $v_{0}$ is a vertex, it is an element of $C$ by definition, so $v_{0}\in\Sigma(\pi)$ and $|\Sigma(\pi)|\geq 1$. Now suppose the length of $\pi$ is 1, say $\pi=v_{0}\rightarrow v_{1}$. This means that $v_{0}\hskip-0.85358pt\!\upharpoonright_{[2,\hskip 1.99168pt{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}]}\cong v_{1}\hskip-0.85358pt\!\upharpoonright_{[1,\hskip 1.99168pt{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}-1]}$, and so by the definition of validity there exists a structure $v$ on ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}+1$ points such that $v\hskip-0.85358pt\!\upharpoonright_{[1,\hskip 1.99168pt{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}]}\cong v_{0}$ and $v\hskip-0.85358pt\!\upharpoonright_{[2,\hskip 1.99168pt{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}+1]}\cong v_{1}$. This structure $v$ has associated path $\pi$, meaning that $|\Sigma(\pi)|\geq 1$. To see that $v\in C$, start by assuming for a contradiction that $\sigma\leq v$ for some $\sigma\in B$. Since $\sigma\nleq v_{0},v_{1}$, it follows that the embedding of $\sigma$ in $v$ must include both the first point of $v_{0}$ and the last point of $v_{1}$. Therefore, $\sigma=v$, since the embedding is contiguous. However, $|\sigma|=|v|=b+1>b$, so $\sigma\notin B$, a contradiction.
Now suppose all paths of length <k<k in $\Gamma_{C}^{m}$ have at least one associated structure. Let $\pi=v_{0}\rightarrow\dots\rightarrow v_{k}$ be a path of length $k$ in $\Gamma_{C}^{m}$. By assumption the paths $v_{0}\rightarrow\dots\rightarrow v_{k-1}$ and $v_{k-1}\rightarrow v_{k}$ each have at least one associated structure; call these $\rho$ and $\sigma$ respectively. Now $\rho$ and $\sigma$ overlap on the $m$ points of $v_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k-1}}$, and so since the structures of this poset are valid there exists a structure $\theta$ which combines $\sigma$ and $\rho$. It is clear that $\Pi(\theta)=\pi$, so $|\Sigma(\pi)|\geq 1$. That $\theta\in C$ can be shown using the reasoning given for the base case.
@@
We will show that forests are not a valid type of structures. Consider the following two forests:
$G\hskip 2.13394pt=132$$H\hskip 2.13394pt=$
It can be seen that $G\hskip-0.85358pt\!\upharpoonright_{[2,3]}\cong H\hskip-0.85358pt\!\upharpoonright_{[1,2]}$. However, any graph which combines $G$ and $H$ contains the cycle shown in Figure 4 as a subgraph, and so is certainly not a forest. Hence forests are not valid structures.
We will now consider how this affects factor graphs. Consider the factor graph in Figure 3, which is the factor graph of the avoidance set $C$ of $\mathcal{F}$ which avoids all forests on three points except for $G$ and $H$. One graph associated with the path $\pi=G\rightarrow H$ is that shown in Figure 4. Since this is a cycle, it is not a forest, and therefore not in $C$. Every graph associated with $\pi$ contains this cycle as a subgraph, so $\pi$ has no associated forests. This means that $\Sigma(\pi)=\varnothing$ and so we can also see that $\mathcal{F}$ is not valid using the contrapositive of Lemma 5.4.
@@
Figure 4. A graph corresponding to the path $\pi$ in Example 5.5
@@
We note that structures which can be defined as downward closed subsets of valid structures are not necessarily valid themselves. One example of this is given by forests: even though they are a downward closed subset of a valid type of structure (namely, graphs), we have seen in Example 5.5 that they are not valid. Another example is given by overlap free words. A word over an alphabet $A$ is overlap free if it does not contain any subword of the form xyxyxxyxyx for $x\in A,y\in A\cup\{\epsilon\}$. Note that this definition is different to the one commonly used in the literature (eg. [1]), where $y$ can be a word. To see that overlap free words are not valid structures, consider the set of overlap free words over $A=\{a,b\}$. Take $u=aba$, $v=baba$; clearly $u,v\in A^{*}$. Moreover, $u\hskip-0.85358pt\!\upharpoonright_{[2,3]}=ba=v\hskip-0.85358pt\!\upharpoonright_{[1,2]}$. However, identifying these two letters yields the word ababaababa, which is not overlap free.
In this case we can use the methodology for words under the consecutive order to answer the wqo and atomicity problems for overlap free words under the consecutive order. Let $\mathcal{OA}_{A}$ be the poset of overlap free words over an alphabet A under the consecutive order. We consider an avoidance set $C=\operatorname{Av}(B)$ of $\mathcal{OA}_{A}$. Take $Y=\{xyxyx\hskip 5.69054pt|\hskip 5.69054ptx\in A,y\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A\cup\{\epsilon\}}\}$. It can be seen that $C$ is equal to the avoidance set $\operatorname{Av}(B\cup Y)$ of $\mathcal{W}_{A}$. Hence wqo and atomicity for $C$ are governed by wqo and atomicity respectively for this avoidance set of $\mathcal{W}_{A}$.
@@ 6. Bountiful structures
A valid type of structure is said to be bountiful if for any structures $\sigma,\rho$ of this type which satisfy
  • $|\sigma|=|\rho|=m$ for some $m\in\mathbb{N}$;
  • $\sigma\hskip-0.85358pt\!\upharpoonright_{[2,m]}\cong\rho\hskip-0.85358pt\!\upharpoonright_{[1,m-1]}$;
there exist at least two structures $\theta_{1},\theta_{2}$ of the same type and of length $m+1$, such that $\theta_{i}\hskip-0.85358pt\!\upharpoonright_{[1,\hskip 1.99168ptm]}\cong\sigma$ and $\theta_{i}\hskip-0.85358pt\!\upharpoonright_{[2,\hskip 1.99168ptm+1]}\cong\rho$ for $i=1,2$.
As for valid structures, if the type of structure of the underlying set of a poset is bountiful, we will abuse the terminology by referring to structures of this type as bountiful structures.
@@
We begin by showing that $(1)$ holds. Given any path $\pi=\sigma\rightarrow\rho$ in $\Gamma^{m}_{C}$, $\sigma\hskip-0.85358pt\!\upharpoonright_{[2,\hskip 1.99168ptm]}\cong\rho\hskip-0.85358pt\!\upharpoonright_{[1,\hskip 1.99168ptm-1]}$ so there are two structures $\theta_{1},\theta_{2}$ on $m+1$ points such that $\theta_{i}\hskip-0.85358pt\!\upharpoonright_{[1,\hskip 1.99168ptm]}\cong\sigma$ and $\theta_{i}\hskip-0.85358pt\!\upharpoonright_{[2,\hskip 1.99168ptm+1]}\cong\rho$, for $i=1,2$, by the assumption of bountifulness. That $\theta_{1},\theta_{2}\in C$ follows by the same argument used in the proof of Lemma 5.4. Both $\theta_{1}$ and $\theta_{2}$ are in $\Sigma(\pi)$, so $|\Sigma(\pi)|\geq 2$ as required.
We will use the fact that $(1)$ holds to prove $(2)$. Let $\pi=v_{1}\rightarrow\dots\rightarrow v_{n}$ be a path in $\Gamma_{C}^{m}$. If $|\pi|=1$ then $\pi$ is ambiguous by $(1)$, so suppose that $|\pi|>1$ (meaning that $n\geq 3$). By $(1)$, the subpath $\eta=v_{1}\rightarrow v_{2}$ is ambiguous, so there exist two distinct structures $\theta_{1},\theta_{2}\in\Sigma(\eta)$. Consider the subpath $\pi^{-}=v_{2}\rightarrow\dots\rightarrow v_{n}$ of $\pi$. By Lemma 5.4, there exists $\sigma\in\Sigma(\pi^{-})$. Now $\sigma$ overlaps with both $\theta_{1}$ and $\theta_{2}$ on the $m$ vertices of $v_{2}$, so since bountiful types of structures are also valid, there exist structures $\rho_{1},\rho_{2}$ which combine $\sigma$ and $\theta_{1},\theta_{2}$ respectively. Note that $\rho_{1}$ and $\rho_{2}$ are distinct because $\theta_{1}$ and $\theta_{2}$ are distinct. Again, that $\rho_{1},\rho_{2}\in C$ follows in the same way as in the proof of Lemma 5.4. To finish, we observe that, since $\Pi(\rho_{1})=\Pi(\rho_{2})=\pi$, it follows that $\rho_{1},\rho_{2}\in\Sigma(\pi)$, and so $\pi$ is ambiguous.
@@
[…]